INFRARED SPECTRUM OF THE CS_{2} TETRAMER: OBSERVATION OF A STRUCTURE WITH D_{2d} SYM-METRY

M. REZAEI, \underline{J. NOROOZ OLIAEE}, N. MOAZZEN-AHMADI, Department of Physics and Astronomy, University of Calgary, 2500 University Dr., N.W., Calgary, Alberta T2N 1N4, Canada; A.R.W. McKELLAR, National Research Council of Canada, Ottawa, Ontario, K1A 0R6, Canada.

Infrared spectra of carbon disulphide clusters are studied in the region of the CS_{2} ν_{3} fundamental band (∼1535 cm^{−1}), using a tuneable diode laser to probe a pulsed supersonic slit jet expansion. A symmetric rotor parallel band at 1551.438 cm^{−1} is assigned to CS_{2} tetramer. The likely structure is a staggered and tilted barrel with D_{2d} symmetry, similar to the previously observed oblate tetramer of nitrous oxide, in reasonable agreement with that calculated from an empirical CS_{2} intermolecular potential function. This is the first high-resolution spectroscopic detection of CS_{2} tetramer, and the calculations suggest that there could be other lower energy isomers of (CS_{2})_{4} which are not yet observed.